xx TestLoopControl

TestLoopControl

Erik Schnetter <schnetter@gmail.com>

December 22 2012

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

7 Interfaces

General Implements: testloopcontrol

Inherits: grid

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  csums  10  x01  compact110
11                 10  csumxall  x01 description11Result values for C test cases
11                 10  csumxint  x01 dimensions110
11                 10  csumxbnd  x01 distribution11CONSTANT
11                 10  csumxintbnd  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11  fsums  10  x01  compact110
11                 10  fsumxall  x01 description11Result values for Fortran test cases
11                 10  fsumxint  x01 dimensions110
11                 10  fsumxbnd  x01 distribution11CONSTANT
11                 10  fsumxintbnd  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 pointtypes 10 pointtypes x01 compact110
11                 10  x01 description11Grid function point types
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels111
11                 10  x01variable type11INT

11

Uses header: loopcontrol.h

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/TestLoopControl. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11csums fsums pointtypes11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x testloopcontrolpointwise x test loop iterators pointwise x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxINITIAL x testloopcontrolc x test loop iterators x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxINITIAL x testloopcontrolfortran x test loop iterators

11x11Language:11fortran
11 x 11Type: 11function